An angle measures 61.8 more than the measurement of its complementary angle. What is the measure of each angle

Found 2 solutions by MathLover1, ikleyn:
Answer by MathLover1(20850)


two angles are called complementary if their measures add to 90 degrees

so if an angle measures 61.8 more than the measurement of its complementary angle, then the measurement of its complementary is90-61.8=28.2 degrees

Answer by ikleyn(52832)

.
An angle measures 61.8 more than the measurement of its complementary angle.
What is the measure of each angle
~~~~~~~~~~~~~~~~~~~~~~~~~~

Let the unknown angle be x degrees.


Then the complementary angle is  (90-x) degrees.


The problem says


    x - (90-x) = 61.8 degrees.


Simplify, solve and find x


    2x = 61.8 + 90

    2x = 151.8

     x = 151.8/2 = 75.9  degrees.    ANSWER

Solved, and thoroughly explained.